# Spin defects in hexagonal boron nitride as two-dimensional strain sensors

### v1

@@ Abstract
Lattice deformation is a powerful way to engineer the properties of two-dimensional (2D) materials, making their precise measurement an important challenge for both fundamental science and technological applications. Here, we demonstrate that boron-vacancy (V) color centers in hexagonal boron nitride (hBN) enable quantitative strain sensing with sub-micrometer spatial resolution. Using this approach, we precisely quantify the strain-induced shift of the E2g Raman mode in a hBN flake under uniaxial stress, establishing V centers as a new tool for strain metrology in van der Waals heterostructures. Beyond strain sensing, our work also highlights the unique multimodal sensing functionalities offered by V centers, which will be valuable for future studies of strain-engineered 2D materials.
@@
In this Letter, we explore a novel approach to strain sensing in 2D materials that relies on a quantum sensing platform based on boron-vacancy (V) color centers in hexagonal boron nitride (hBN) Vaidya et al. (2023). This photoluminescent point defect features a spin-triplet ground state, whose electron spin resonance (ESR) frequencies can be interrogated via optically detected magnetic resonance (ODMR) spectroscopy under ambient conditions Gottscholl et al. (2020), even in the limit of ultra-thin hBN layers Durand et al. (2023). In addition, the ESR frequencies are highly sensitive to external perturbations such as magnetic and electric fields, pressure, temperature, or strain Liu et al. (2021); Gottscholl et al. (2021); Udvarhelyi et al. (2023); Yang et al. (2022); Lyu et al. (2022). These properties make V centers in hBN a promising candidate for the development of flexible, multifunctional sensors that can be readily integrated into van der Waals heterostructures for probing the physics of 2D materials with atomic-scale proximity Huang et al. (2022); Healey et al. (2023); Kumar et al. (2022); Mu et al. (2025); He et al. (2025). Here, we present the first experimental study of V centers under continuously tunable in-plane strain, exploiting the flexibility of thin hBN flakes as a host. Using an ensemble of V centers, we demonstrate quantitative strain sensing with sub-micrometer spatial resolution. We then use this capability to precisely calibrate the strain induced in a hBN flake under uniaxial stress, and we investigate the strain-induced shift of the E2g Raman mode. Despite the central role of hBN in 2D materials research, the impact of strain on its vibrational properties has received surprisingly little attention, with only a few studies published in the literature Cai et al. (2017); Androulidakis et al. (2018); Blundo et al. (2022). Our results provide a precise quantification of the strain dependence of the E2g Raman mode in hBN, revealing a shift of cm, which can serve as a reference for future studies of strain in van der Waals heterostructures. Beyond this result, our work highlights the potential of V centers as a robust tool for multimodal sensing in van der Waals heterostructures, with important implications for the study of strain-engineered 2D materials.
@@
where are the dimensionless electron spin operators. The first term results from the spin-spin interaction between the two unpaired electrons of the defect, which leads to an axial zero-field splitting GHz between the and spin sublevels Gottscholl et al. (2020), where denotes the electron spin projection along the axis () of the hBN crystal. The second term arises from the interaction of the V center with a local electric field produced by surrounding charges in the hBN flake Gong et al. (2023); Durand et al. (2023); Udvarhelyi et al. (2023), which mixes the spin sublevels, yielding new eigenstates separated by an orthorhombic splitting [Fig. 2(a)]. This splitting increases with the density of charges in the hBN crystal and thus with the density of V centers. For the hBN flakes used in this work, MHz. The last term finally describes the spin-strain interaction, which is expressed as Udvarhelyi et al. (2023)
@@
First-principles calculations predict a strain coupling coefficient of MHz/% for a hBN monolayer, which increases to MHz/% for V centers embedded in thicker hBN flakes such as those used in this work Udvarhelyi et al. (2023). Tensile strain (taken positive by convention) thus leads to a reduction of the effective zero-field splitting parameter [Fig. 2(a)]. These theoretical predictions have been quantitatively confirmed by several recent experiments performed under hydrostatic stress conditions Mu et al. (2025); He et al. (2025); Yu et al. (2025). In the following, we employ the theoretically predicted coefficient to extract the local strain in the hBN flake from measurements of the ESR frequencies.
@@
V spin sensors can also be used to infer the strain distribution within the hBN flake with a diffraction-limited spatial resolution (m). Figure 3 shows typical strain profiles recorded along the long axis () of the hBN flake for various voltages applied to the strain cell. These measurements reveal a slight strain gradient of about %/m, which emerges between the two sides of the strain cell. This gradient likely results from different adhesion of the hBN flake on the strain-transmitting substrate at the opposite edges of the gap
@@
As a first application of V-based strain sensing, we now investigate the strain-induced variation of the E2g Raman mode in hBN, a topic scarcely documented in the literature despite the central role of hBN in 2D materials research. To this end, both Raman and ODMR spectra are sequentially measured at the center of the hBN flake while gradually increasing the voltage applied to the strain cell. The ODMR spectra provide a precise, in situ strain calibration at the exact location where Raman spectra are recorded. As shown in Fig. 4(a), the effective zero-field splitting parameter of the V center initially decreases linearly with the applied voltage, consistently with the previous experiment. It reaches a minimum value GHz at a voltage of V, corresponding to a maximum strain of . Interestingly, further elongation of the strain cell does not increase the strain in the hBN flake, but instead leads to its gradual release as the voltage rises from V to V, most likely due to layer slippage Androulidakis et al. (2018). When the voltage is swept back to V, the strain decreases linearly with a slope similar to that observed during the voltage ramp-up and reaches a compressive state at zero voltage [Fig. 4(a)]. This pronounced hysteretic behavior of strain in a hBN flake under uniaxial stress highlights the complexity of the mechanical response in layered 2D materials, which cannot be captured by standard numerical simulations.

### v2

@@ Abstract
Lattice deformation is a powerful way to engineer the properties of two-dimensional (2D) materials, making their precise measurement an important challenge for both fundamental science and technological applications. Here, we demonstrate that boron-vacancy (V) color centers in hexagonal boron nitride (hBN) enable quantitative strain sensing with sub-micrometer spatial resolution. Using this approach, we precisely quantify the strain-induced shift of the E2g Raman mode in a multilayer hBN flake under uniaxial stress, establishing V centers as a new tool for strain metrology in van der Waals heterostructures. Beyond strain sensing, our work also highlights the unique multimodal sensing functionalities offered by V centers, which will be valuable for future studies of strain-engineered 2D materials.
@@
In this Letter, we explore a novel approach to strain sensing in 2D materials that relies on a quantum sensing platform based on boron-vacancy (V) color centers in hexagonal boron nitride (hBN) Vaidya et al. (2023). This photoluminescent point defect features a spin-triplet ground state, whose electron spin resonance (ESR) frequencies can be interrogated via optically detected magnetic resonance (ODMR) spectroscopy under ambient conditions Gottscholl et al. (2020), even in the limit of ultra-thin hBN layers Durand et al. (2023). In addition, the ESR frequencies are highly sensitive to external perturbations such as magnetic and electric fields, pressure, temperature, or strain Liu et al. (2021); Gottscholl et al. (2021); Lyu et al. (2022); Udvarhelyi et al. (2023); Yang et al. (2022). These properties make V centers in hBN a promising candidate for the development of flexible, multifunctional sensors that can be readily integrated into van der Waals heterostructures for probing the physics of 2D materials with atomic-scale proximity Huang et al. (2022); Healey et al. (2023); Kumar et al. (2022); Mu et al. (2025); He et al. (2025). Here, we present the first experimental study of V centers under continuously tunable in-plane strain, exploiting the flexibility of thin hBN flakes as a host. Using an ensemble of V centers, we demonstrate quantitative strain sensing with sub-micrometer spatial resolution. We then use this capability to precisely calibrate the strain induced in a multilayer hBN flake under uniaxial stress, and we investigate the strain-induced shift of the E2g Raman mode. Despite the central role of hBN in 2D materials research, the impact of strain on its vibrational properties has received surprisingly little attention, with only a few studies published in the literature Cai et al. (2017); Androulidakis et al. (2018); Blundo et al. (2022). Our results provide a precise quantification of the strain dependence of the E2g Raman mode in multilayer hBN, revealing a shift of cm, which can serve as a reference for future studies of strain in van der Waals heterostructures. Beyond this result, our work highlights the potential of V centers as a robust tool for multimodal sensing in van der Waals heterostructures, with important implications for the study of strain-engineered 2D materials.
@@
where are the dimensionless electron spin operators. The first term results from the spin-spin interaction between the two unpaired electrons of the defect, which leads to an axial zero-field splitting GHz between the and spin sublevels Gottscholl et al. (2020), where denotes the electron spin projection along the axis () of the hBN crystal. The second term arises from the interaction of the V center with a local electric field produced by surrounding charges in the hBN flake Gong et al. (2023); Durand et al. (2023); Udvarhelyi et al. (2023), which mixes the spin sublevels, yielding new eigenstates separated by an orthorhombic splitting [Fig. 2(a)]. This splitting increases with the density of charges in the hBN crystal and thus with the density of V centers. For the hBN flakes used in this work, MHz, corresponding to a density of V centers around ppm Gong et al. (2023); Udvarhelyi et al. (2023). The last term finally describes the spin-strain interaction, which is expressed as Udvarhelyi et al. (2023)
@@
First-principles calculations predict a strain coupling coefficient of MHz/% for a hBN monolayer, which increases to MHz/% for V centers embedded in multilayer hBN flakes such as those used in this work Udvarhelyi et al. (2023). Tensile strain (taken positive by convention) thus leads to a reduction of the effective zero-field splitting parameter [Fig. 2(a)]. These theoretical predictions have been quantitatively confirmed by several recent experiments performed on multilayer hBN flakes under hydrostatic stress conditions Mu et al. (2025); He et al. (2025); Yu et al. (2025). We note that the strain coupling coefficient is one order of magnitude smaller than that reported in earlier works Gottscholl et al. (2021); Lyu et al. (2022). In those studies, the strain susceptibility of the V center was indirectly inferred from the temperature dependence of the axial zero-field splitting parameter, assuming it arises solely from an effective stress due to thermal distortion of the hBN lattice. More recent work has shown that the variations of the zero-field splitting parameter with temperature instead originate from temperature-dependent spin-phonon interactions Liu et al. (2025), explaining the discrepancy in the reported strain susceptibilities. In the following, we employ the theoretically predicted and experimentally confirmed strain coupling coefficient to extract the local strain in a multilayer hBN flake from measurements of the ESR frequencies.
@@
V spin sensors can also be used to infer the strain distribution within the hBN flake with a diffraction-limited spatial resolution (m). Figure 3 shows typical strain profiles recorded along the long axis () of the hBN flake for various voltages applied to the strain cell. These measurements reveal a slight strain gradient of about %/m, which emerges between the two sides of the strain cell. This gradient likely results from different adhesion of the hBN flake on the strain-transmitting substrate at the opposite edges of the gap.
@@
As a first application of V-based strain sensing, we now investigate the strain-induced variation of the E2g Raman mode in hBN, a topic that remains scarcely documented in the literature despite the central role of hBN in 2D materials research. To this end, both Raman and ODMR spectra are sequentially measured at the center of the hBN flake while gradually increasing the voltage applied to the strain cell. The ODMR spectra provide a precise, in situ strain calibration at the exact location where Raman spectra are recorded. As shown in Fig. 4(a), the effective zero-field splitting parameter of the V center initially decreases linearly with the applied voltage, consistently with the previous experiment. It reaches a minimum value GHz at a voltage of V, corresponding to a maximum strain of . Interestingly, further elongation of the strain cell does not increase the strain in the hBN flake, but instead leads to its gradual release as the voltage rises from V to V. This effect most likely result from slippage between the multilayer hBN flake and the PI substrate Androulidakis et al. (2018). We note that the ODMR linewidth is not modified in this regime, suggesting that in-plane strain does not vary significantlly across the thickness of the hBN flake. When the voltage is swept back to V, the strain decreases linearly with a slope similar to that observed during the voltage ramp-up and reaches a compressive state at zero voltage [Fig. 4(a)]. This pronounced hysteretic behavior of strain in a hBN flake under uniaxial stress highlights the complexity of the mechanical response in layered 2D materials, which cannot be captured by standard numerical simulations.